OpenStudy (marissalovescats):

@jim_thompson5910 Alright let's finish this Part 1 We left off on 89

jimthompson5910 (jim_thompson5910):

yeah that sounds right

jimthompson5910 (jim_thompson5910):

so tell me what you have for 89

OpenStudy (marissalovescats):

What I have is nada

OpenStudy (marissalovescats):

Zilch

jimthompson5910 (jim_thompson5910):

g(x) = x*f(x) that's given

jimthompson5910 (jim_thompson5910):

we want to find the tangent off g(x) so we need g ' (x) to get the slope of the tangent line

OpenStudy (marissalovescats):

Yis

jimthompson5910 (jim_thompson5910):

when you derive x*f(x), you have to use the product rule

jimthompson5910 (jim_thompson5910):

so go ahead and do that for me and tell me what you get

OpenStudy (marissalovescats):

Okay

OpenStudy (marissalovescats):

\[g'(x) = f(x)+x*f'(x)\]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now evaluate g ' (2) this yields the slope of the tangent line at x = 2 (on g(x))

OpenStudy (marissalovescats):

g'(2)=-7

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now you need g(2)

jimthompson5910 (jim_thompson5910):

this is so you know the point (x,y) that lies on g(x) when x = 2

OpenStudy (marissalovescats):

y-6=-7(x-2) so D

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (marissalovescats):

Hm no idea about 90

jimthompson5910 (jim_thompson5910):

f has a positive first derivative

jimthompson5910 (jim_thompson5910):

what does that mean?

OpenStudy (marissalovescats):

The slope is always positive

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (marissalovescats):

It's above the x axis?

jimthompson5910 (jim_thompson5910):

f ', yes

jimthompson5910 (jim_thompson5910):

what about increasing or decreasing?

OpenStudy (marissalovescats):

It's always increasing

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so which choices can you eliminate based on this fact alone?

OpenStudy (marissalovescats):

C, D and E

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

then we move onto the fact that f '' (x) < 0 for all x ie a negative second derivative

jimthompson5910 (jim_thompson5910):

do you remember how the slopes of the secant lines correspond to concavity?

OpenStudy (marissalovescats):

I don't even think we talked about secant lines

OpenStudy (marissalovescats):

Or I just don't remember.

jimthompson5910 (jim_thompson5910):

here's some concave up curve 

OpenStudy (marissalovescats):

Yes

jimthompson5910 (jim_thompson5910):

and some points that lie on it